A provider order is written for insulin at 3 units/hr; the premixed solution is composed of 50 units of insulin in 100 ml in normal saline. At which rate (ml/hr) should the infusion run?
Answer #1

ANSWER.

* Calculation of rate(ml/hr) should the insulin run.

            Formula

                Flow rate = Volume(ml) /Time(hr)

      Volume(ml) of insulin required.

              Formula = Desired dose(D) / Dose in hand(H) * Quantity(Q)

                    Here

                      Desired dose(D) = 3 units/hr.

                      Dose in hand (H) = 50 units.

                           Quantity(Q) = 100 ml.

    As per formula

           Volume(ml) of insulin required = D/H * Q

                                                            = 3/50 * 100

                                                            = 0.06 * 100

                                                            = 6 ml.

                                        Time(hr) = 1 hr.

                    Flow rate (ml/hr) = Volume(ml) / Time(hr)

                                                          = 6/1

                                                          = 6 ml/hr.